# Open foundation models for Azerbaijani language

### v1

@@ 2 Previous works
There have also been some community attempts to create such open-source models. A series of RoBERTa models were developed by continuing the pre-training phase on a small Azerbaijani dataset (Hajili, 2024c). Alas Development Center has developed a series of decoder-only LLMs for Azerbaijani  [^3] https://github.com/interneuron-ai/project-barbarossa, but they offer no explanation regarding their approach, and the models failed to pass initial sanity checks.
@@ 4 Pre-training
Using DOLLMA, we have developed a series of foundational language models called aLLMA (a Large Language Model for Azerbaijani). aLLMA has been trained in three sizes: small, base, and large. Base and large correspond to the original BERT models BERTBASE and BERTLARGE (Devlin et al., 2019). Small architecture was borrowed from (Bhargava et al., 2021). Architectural details of these models can be found in Table 2. aLLMA-SMALL [^9] https://huggingface.co/allmalab/bert-small-aze and aLLMA-BASE [^10] https://huggingface.co/allmalab/bert-base-aze have been trained and are included in our benchmarks. aLLMA-LARGE will be released before September, 2024 and the benchmarks will be updated accordingly.
@@
SQuAD (Translated) 54.1k Extractive QA (Hajili, 2024d)
@@
Figure 1: Training loss for aLLMA-SMALL aLLMA-BASE and aLLMA-LARGE models.
@@ 5.4.2 SQuAD
Question-answering problems usually demand more robust language understanding and therefore serve as a better criterion than simpler classification tasks. There is no original open-book question-answering dataset in Azerbaijani. The Stanford Question Answering Dataset (SQuAD) is one such dataset in English. We used a translated and reindexed version of the original (Hajili, 2024d).
@@ 6 Results
Model name Size AZE-SCI MRPC WikiANN SQuAD AZE-NSP LDQuAd
XLM-RoBERTa-LARGE 560M 89.76 82.41 92.35 75.70 33.46 83.48
mDeBERTa-v3-BASE 279M 87.13 83.71 91.87 72.27 78.84 85.29
XLM-RoBERTa-BASE 278M 86.99 70.90 90.29 70.97 74.96 85.17
RoBERTa-BASE-AZE 278M 89.17 81.25 91.62 70.36 76.98 85.44
BERT-BASE-AZE 178M 88.80 80.12 92.35 69.42 74.12 64.41
BERT-BASE-MULTI 178M 86.88 79.92 91.67 68.92 72.46 83.48
BERT-SCRATCH 135M 73.31 65.36 72.95 16.11 50.73 26.60
BERT-BASE 108M 76.73 75.00 90.94 55.51 62.12 74,88
ALLMA-BASE 135M 90.84 79.74 91.26 71.30 75.95 86.26
ALLMA-SMALL 46M 88.06 71.77 90.07 59.89 70.23 80.80
@@
Multilingual foundation models. BERT-BASE-MULTI is a multilingual version of the original BERT model. XLM-RoBERTa-BASE and XLM-RoBERTa-LARGE are some of the best-performing multilingual models (Conneau et al., 2020). mDeBERTa-v3-BASE is a multilingual version of DeBERTa v3 model (He et al., 2023)).
Multilingual models further pre-trained for Azerbaijani. BERT-BASE-AZE (Hajili, 2024b) and RoBERTa-BASE-AZE (Hajili, 2024c) have been further pre-trained on a small and high-quality Azerbaijani dataset. Their base models are RoBERTA-BASE, BERT-BASE-MULTI, and DeBERTa-BASE, respectively.
Models pre-trained from scratch. aLLMA-SMALL and aLLMA-BASE are the only monolingual Azerbaijani models. aLLMA-LARGE is still being trained.
@@
You can find the results in Table 4. mDeBERTa-v3-BASE and aLLMA-BASE have the best overall performance. Figure 2 compares the performance of BASE models. [^12] The difference in number of parameters between these models is due to varying vocabulary sizes. Otherwise, their architectures are identical. aLLMA-BASE outperforms all other models of similar size in 4 out of 6 benchmarks. Comparing BERT-BASE-AZE with BERT-BASE-MULTI shows that further pre-training of multilingual models can result in some performance improvement, but also model collapse (compare their performance in LDQuAd benchmark). However, a more comprehensive analysis is required before we can make generalizations about the effects of continued monolingual pre-training on multilingual models.
@@
    Multilingual models offer competitive performance even in languages that they were undertrained for. Azerbaijani has not been the focus in any of these multilingual models (XLM-RoBERTa, mDeBERTa-v3-BASE, or BERT-BASE-MULTI). Despite this, they outperform most models in some tasks.
@@ Acknowledgements
This work has been funded by PRODATA LLC and performed at aLLMA Lab 11 . We would like to thank Haim Dror and Mirakram Aghalarov for the invaluable discussions that led to this project. We would also like to thank the BHOS AI team for their close collaboration throughout the project. All of the data and code that has been created as part of this project is going to be publicly available under permissive licenses.

### v2

@@ 2 Previous works
There have also been some community attempts to create such open-source models. A series of RoBERTa models were developed by continuing the pre-training phase on a small Azerbaijani dataset (Hajili, 2024d). Alas Development Center has developed a series of decoder-only LLMs for Azerbaijani  [^3] https://github.com/interneuron-ai/project-barbarossa, but they offer no explanation regarding their approach, and the models failed to pass initial sanity checks.
@@ 4 Pre-training
Using DOLLMA, we have developed a series of foundational language models called aLLMA (a Large Language Model for Azerbaijani). aLLMA has been trained in three sizes: small, base, and large. Base and large correspond to the original BERT models BERTBASE and BERTLARGE (Devlin et al., 2019). Small architecture was borrowed from (Bhargava et al., 2021). Architectural details of these models can be found in Table 2. All three models [^9] https://huggingface.co/allmalab/bert-small-aze, [^10] https://huggingface.co/allmalab/bert-base-aze, [^11] https://huggingface.co/allmalab/bert-large-aze have been released publicly and included in our benchmarks.
@@
SQuAD (Translated) 54.1k Extractive QA (Hajili, 2024e)
@@
Figure 1: Training loss for aLLMA-Small, aLLMA-Base, and aLLMA-Large models.
@@ 5.4.2 SQuAD
Question-answering problems usually demand more robust language understanding and therefore serve as a better criterion than simpler classification tasks. There is no original open-book question-answering dataset in Azerbaijani. The Stanford Question Answering Dataset (SQuAD) is one such dataset in English. We used a translated and reindexed version of the original (Hajili, 2024e).
@@ 6 Results
Model name Size AZE-SCI MRPC WikiANN SQuAD AZE-NSP LDQuAd Avg.
XLM-RoBERTa-Large 560M 89.76 82.41 92.35 75.70 33.46 83.48 76.19
mDeBERTa-v3 279M 87.13 83.71 91.87 72.27 78.84 85.29 83.19
mDEBERTA-v3-AZE 279M 89.73 80.18 91.83 70.31 78.29 85.07 82.57
XLM-RoBERTa-Base 278M 86.99 70.90 90.29 70.97 74.96 85.17 79.88
RoBERTa-Base-AZE 278M 89.17 81.25 91.62 70.36 76.98 85.44 82.47
BERT-Base-AZE 178M 88.80 80.12 92.35 69.42 74.12 64.41 78.20
BERT-Base-Multi 178M 86.88 79.92 91.67 68.92 72.46 83.48 80.56
BERT-Scratch 135M 73.31 65.36 72.95 16.11 50.73 26.60 50.84
BERT-Base 108M 76.73 75.00 90.94 55.51 62.12 74.88 72.53
ALLMA-Large 370M 91.46 81.55 91.71 73.77 78.58 85.93 83.83
ALLMA-Base 135M 90.84 79.74 91.26 71.30 75.95 85.69 82.46
ALLMA-Small 46M 88.06 71.77 90.07 59.89 70.23 80.80 76.80
@@
Multilingual foundation models. BERT-Base-MULTI is a multilingual version of the original BERT model. XLM-RoBERTa-Base and XLM-RoBERTa-Large are some of the best-performing multilingual models (Conneau et al., 2020). mDeBERTa-v3 is a multilingual version of DeBERTa v3 model (He et al., 2023)).
Multilingual models further pre-trained for Azerbaijani. BERT-Base-AZE (Hajili, 2024b), RoBERTa-Base-AZE (Hajili, 2024d), and mDEBERTA-v3-AZE (Hajili, 2024c) have been further pre-trained on a small and high-quality Azerbaijani dataset. Their base models are RoBERTA-Base, BERT-Base-MULTI, and DeBERTa-Base, respectively.
Models pre-trained from scratch. aLLMA-Small, aLLMA-Base, and aLLMA-Large are the only monolingual Azerbaijani models.
@@
You can find the results in Table 4. mDeBERTa-v3 and aLLMA-Base have the best overall performance. Figure 2 compares the performance of Base models. [^13] The difference in number of parameters between these models is due to varying vocabulary sizes. Otherwise, their architectures are identical. aLLMA-Base outperforms all other models of similar size in 4 out of 6 benchmarks. Comparing BERT-Base-AZE with BERT-Base-MULTI shows that further pre-training of multilingual models can result in some performance improvement, but also model collapse (compare their performance in LDQuAd benchmark). However, a more comprehensive analysis is required before we can make generalizations about the effects of continued monolingual pre-training on multilingual models.
@@
    Multilingual models offer competitive performance even in languages that they were undertrained for. Azerbaijani has not been the focus in any of these multilingual models (XLM-RoBERTa, mDeBERTa-v3, or BERT-Base-MULTI). Despite this, they outperform most models in some tasks.
@@ Acknowledgements
We thank PRODATA LLC and Microsoft Accelerating Foundation Models Research Program for supporting our research. All of the data and code created as part of this project will be publicly available under permissive licenses.